OpenStudy (anonymous):

Use function notation to write the equation of the line. bit.ly/1ffoCNs A. f (x) = 2x – 1 B. f (x) = –2x – 1 C. f (x) = –½x – 1 D. f (x) = –2x + 1

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

f(x)=mx+b where m is the slope, and b is the y-intercept

OpenStudy (anonymous):

But, I dont even know what to plug into the unknown variables

OpenStudy (unklerhaukus):

look at the plot, what value does the line cross the y-axis? this is the y-intercept b

OpenStudy (anonymous):

(0,-1)?

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

So b=-1?

OpenStudy (unklerhaukus):

yes

OpenStudy (anonymous):

f(x)=mx+(-1)

OpenStudy (unklerhaukus):

that is the y-intercept so your almost there

OpenStudy (unklerhaukus):

now you just need to get the slope (or gradient )

OpenStudy (unklerhaukus):

the gradient is the amount the height changes , when the vertical displacement is positive one unit

OpenStudy (unklerhaukus):

or equivalently slope is rise/run between any two points 

OpenStudy (unklerhaukus):

i like to choose a run of +1 to make it easy

OpenStudy (anonymous):

Ok

OpenStudy (unklerhaukus):

m = rise/run = (y_2-y_1)/(x_2-x_1) choose two points, say (x_1, y_1) = (0,-1) and say (x_2, y_2) = (1, -3) (if you want to you can choose any two points)

OpenStudy (anonymous):

Ok, I will just use the points you used

OpenStudy (anonymous):

Y=2 X=-1?